United Millers imports wheat from U.S.A at initial cost of 350 dollar per tone.The
shipping costs and customs duty are then charged as 25% and 15% respectively.
When the wheat reached Mombasa,an 8% of the initial cost is incurred to transport
it to Kisumu.
(a)Given that 1 US dollar = Ksh 82.40 ,calculate the total cost of importing 5 tonnes of
wheat in Ksh.
(b)The united millers intend to make a profit of 25%. Giving your answer to the
nearest ten cents,calculate the price at which a 2 kg packet of wheat should be
sold.
c.How much profit shall the company realize from the sale of 1 tonne of wheat.
1
Expert's answer
2019-11-26T10:06:30-0500

(a) Total cost of importing 1 tonne: "350*(1+0.25+0.15+0.08)=\$518."

Total cost of importing 5 tonnes: "518*5=\$2590."

Total cost in Ksh: "2590*82.40=213416."


(b) The price of 2 kg packet: "\frac{518(1+0.25)}{1000}*2=\$1.30."

The price of 2 kg packet in Ksh: "1.30*82.40=107.10."


(c) Profit from the sale of 1 tonne: "518*0.25=\$129.50."

Profit in Ksh: "129.50*82.40=10670.80."